# Beyond the Carnot limit: work extraction via an entropy battery

### v1

@@ III.1 Distinguishability
where $\mathbf{k}=\{k_{0},k_{1},...,k_{d-1}\}$ specifies a configuration of the $k_{j}$ particles over the $j$ states. See Appendix B.1 for further details. This partition function is immediately factorizable for every one of the NN particles, since it is identically a multinomial expansion
@@
where $p(\gamma)$ is the single particle Boltzmann distribution. The entropy at infinite temperature $\gamma=0$ is therefore proportional to the number of particles
@@
A separable probability distribution and extensive entropy represents mutually independent entities. In the distinguishable particle ensemble, this mutual independence holds for each of the NN particles. Mutual independence therefore implies non-interaction, where interaction is defined generally as an event in which particles exchange a conserved quantity under the constraint of the relevant conservation law, or as a transfer of mutual information between particles. Consequently, distinguishable particles that maintain distinguishability at all times must be non-interacting. This has important implications for the entropy capacity of a realizable entropy battery. For unitary heat engines to operate, or for two canonical ensembles to thermalize, interactions between particles are required. For these reasons, distinguishable particles cannot be used in the entropy battery. Further discussion can be found in Appendix F.
The next best scenario is an ensemble of bosons. As discussed in Appendix B.2, unlike distinguishable particles, the probability distribution for an ensemble of bosons is non-separable down to the individual particles. Therefore, as a contrasting argument to the distinguishable particles, bosons must allow interaction, such that one cannot describe the particles as individuals after there becomes an exchange of information. The inseparability of the probability distributions after interaction is exactly how one would describe particle indistinguishability, and entanglement. The description we give implies that it is the mutual information upon interaction that ultimately determines a particles distinguishability. If the information is mutually exclusive then there remains a form of distinguishability, yet no interaction, but if the information is mutually dependent, then we lose distinguishability and there must have been an interaction. In this sense, what is ultimately indistinguishable upon interactions is the information that is encoded in the conserved quantities of the system. This is especially apparent when monitoring the mutual exchange of information between typically mutually independent conserved quantities like that used in the entropy battery. The exchange of generalized heats between conserved quantities is best understood in the context of information transfer.
@@
IV Conclusion
From the perspective of information thermodynamics, an ensemble of particles with multiple, mutually independent conserved quantities exhibits an expanded entropy capacity. By applying unitary operations that couple these conserved quantities, this entropy capacity can be harnessed during work extraction, yielding energy efficiencies beyond the Carnot limit. We refer to such an ensemble as the entropy battery. We further justify treating spin angular momentum as an independent thermodynamic system by showing that, like energy, it obeys the laws of generalized thermodynamics, follows Bose-Einstein and Fermi-Dirac statistics, satisfies the fluctuation-dissipation and equipartition theorems (see Appendix B), and has analogous Einstein solid and Debye models of heat capacity. We also discuss the conditions under which particles are distinguishable (Appendices B.1 and F), concluding that distinguishable particles can exist only if they are non-interacting. Consequently, the apparent entropy capacity of such particles cannot be used for energetic work extraction via heat engines. This leads naturally to a generalization of indistinguishability in informational terms, independent of the specific conserved quantities or particles, and provides a conceptual understanding of the entropy battery as a device that transfers entropy through subsystem entanglement.
@@ Appendix A Information Thermodynamics
In this section we derive, and present where assumed, the laws of spin thermodynamics, with these laws applying equally to arbitrary conserved quantities.
@@
For ensembles with large NN, the fluctuations of spin angular momentum and entropy between the baths become vanishingly small with respect to the average equilibrium values, as the relative variation scales with $1/\sqrt{N}$ [36]. We can therefore approximate the small changes as an infinitesimal, and make use of the result in eq.(33) to show that this can only be true if
@@ B.1.1 Polarization, spin temperature relation
It is convenient to re-express the spin temperature $\tau$ in terms of the spin-polarization $\alpha$, thus providing a mapping to the macroscopic observable $J_{z}$. From eq.(35)
@@ B.2 Bosons
By changing the index from the particle configuration $\mathbf{k}$ to the macrostate $a(\mathbf{k})=\sum_{j=0}^{d-1}m_{j}k_{j}$, we show that spin bosons follow Bose-Einstein statistics [35]. The macrostate index $a(\mathbf{k})$ ranges from SN-SN to SNSN, which we can shift to the computational macrostate $m=a(\mathbf{k})+SN=\sum_{j=0}^{d-1}jk_{j}$ without affecting the distribution. The multiplicity using this index is derived in Appendix C, and is explicitly the coefficient of the $x^{m}]$th term of a Gaussian binomial,
@@
where $t^{N}]$ is a selection variable, used to select the polynomials of $t$ that correspond to NN particles, and eq.(48) is an identity from Newton’s generalized binomial theorem. Eq.(48) demonstrates that, in the infinite-particle limit, the boson partition function becomes separable over the macrostates
@@ B.3 Fermions
where $t^{N}x^{m}]$ selects the polynomial terms that correspondingly have NN atoms with macrostate index $m$. The fermion partition function using the $m$ index is then
@@
The fermion partition function is thus separable over the macrostates $m$, allowing each state to be treated as a mutually independent grand canonical ensemble, where
@@
which is the Fermi-Dirac distribution [50]. We have removed the selection variable $t$, as this expression does not impose a constraint on the total particle number NN. The separability of the probability distribution across the states $j$ allows each state to be treated as mutually independent. However, as in the bosonic case, the particles themselves cannot be considered mutually independent. Consequently, from a statistical perspective, fermions also permit interaction between particles. Finally, in the limit that $\tau\rightarrow\infty$, eq.(52) becomes
@@ Appendix C Multiplicity of Boson spin states
The number of configurations of indistinguishable bosons that yield the same total spin angular momentum can be reformulated as a combinatorial problem: counting the number of unique paths on an $N\times d$ grid, starting from the bottom-left corner and ending at the top-right corner, with movement restricted to upward and rightward steps, as illustrated in Fig.6. The grid represents NN particles distributed among $d$ states, and the area of the grid above the path taken represents the macrostate $m(\mathbf{k})=\sum_{j=0}^{d-1}jk_{j}$.
@@ Appendix D Waste and Entropic responses
as the amount of entropy the system can store per unit change in spin temperature. It should be clear that any mutually independent conserved quantity possesses an analogous heat capacity. It is therefore natural to define $C_{s}/\tau$ as the entropic response of the ensemble, a quantity that applies generally to any independent conserved variable, such as spin or energy. $C_{s}$ should be interpreted as the waste response of the ensemble, specific to the conserved quantity in question and associated with an analogous “heat” or the loss of usefulness of that conserved quantity. Re-naming $C_{s}$ from what is conventionally the ‘heat’ capacity emphasizes that $C_{s}$ is not a measure of capacity per se, but rather quantifies the absorbed ‘waste’ in response to a change in temperature. This is especially evident when considering the expression of the distinguishable waste response from table 2. The proportionality to the variance of the conserved quantity is a phenomena found generally in the fluctuation-dissipation theorem [36], where $C_{s}$ is the response function. Renaming the heat capacity to the waste response is therefore justified from this perspective. Eqs.(53) and (55) are presented as alternate, yet equivalent interpretations of a systems ability to store waste, both from a physical and an entropic perspective. These two equations allow for us to interpret heat engines in §II as being machines that either filter conserved quantities by separating the ‘heat’ from the ‘work’, or machines that merely transport entropy and information between the heat baths, with these interpretation also applying to the operation of conventional energy heat engine. The information interpretation of the heat engine is general, and leads naturally to concepts involving information transfer between ‘waste’ baths of different conserved quantities, such as spin and energy.
@@
Finally, in the infinite NN limit the entropy and spin therm for bosons becomes eqs.(25) and (2). The boson spin therm is proportional to $\tau$ in the high temperature limit, as the Taylor expansion of $\coth(1/x)\approx x$ for $x\gg 0$, which re-derives the usual heat expression $\mathcal{Q}\propto\tau$ used in, for example, deriving the standard Carnot efficiency limit $\eta=1-\tau_{c}/\tau_{h}$. Eqs.(25) and (2) are presented graphically in Fig.8.
@@ E.1 Distinguishable particles
From eq.(39), the partition function for an ensemble of NN distinguishable particles with $d$ states per particle is
@@
for probability $P(j,\tau)=e^{-j/\tau}/Z$, and $Z=\sum_{j=0}^{d-1}e^{-j/\tau}$. Note also that $\braket{j}N=\braket{\bar{J}_{z}}+SN$. The second derivative is then
@@
$\frac{C_{s}}{N}=\frac{1}{\tau^{2}}\text{Var}(\bar{J}_{z}(\tau))$
@@ E.2 Bosons
From appendix B.2, the partition function for an ensemble of NN boson with $d$ states per particle is
@@ E.3 Debye and Einstein models
This is for 3 degrees of freedom with NN particles, energy $h\nu$, and units J/K thanks to the presence of the Boltzmann constant. For 1 degree of freedom, the unitless, per particle heat capacity is
@@
where we have made the equivalent substitution in the spin degree of freedom, with $2S$ typically being the Debye cut off frequency $\nu_{D}$. Eq.(61) is equal to the boson model (60). The Debye model approximates the sum using an integral, while including the density of frequency modes [37], which is $j^{2}dj$ in the spin basis, resulting in the un-normalized Debye model
@@ Appendix F Distinguishability discussion
With the goal of this paper being to present a realizable entropy storage device, and with entropy capacity translating into extractable energetic work as demonstrated by the entropy battery, then we require an understanding of how one would could increase entropy capacity by means other than enlarging the size of the system. As already discussed in §III and appendix B.1, aside from increasing the number of particles, number of states or the number of independent conserved quantities, the only other apparent means of increasing entropy capacity is to render each particle distinguishable. However, as also established in those sections, a system of distinguishable particles must necessarily be non-interacting. Since particles do in fact interact and heat capacities for distinguishable particles are very different from experimental heat capacities, then particles must be indistinguishable. In this section, we examine whether it is possible to retain the physical entropy capacity associated with distinguishable particles while still allowing interaction. Our conclusion, inspired largely by Saunders’ analysis of indistinguishability [55], is that this is not possible even in principle. The reader is referred to his work for a thorough treatment of the indistinguishability problem.
@@
We begin by clarifying what we mean by distinguishability. From an entropic perspective, physical distinguishability increases the number of accessible microstates $\Omega$ available to the ensemble, leading to an enlarged entropy capacity. The total number of microstates is determined by the number of unique distributions of the particles in state space. Under this definition, if two ensemble states at the same temperature are physically identifiable, then these are distinct microstates that contribute to the system’s physical entropy and hence to its extractable work.
@@
Consider first a single classical particle with physical properties such as position, momentum, and charge. If these properties are localised in state space (a minimum entropy state), then the particle can be tracked continuously through unitary evolution, rendering it distinguishable at all times. Now introduce a second identical particle (same physical properties), however occupying a different location in state space. As long as these particles do not interact and their probability distributions remain separable, they are distinguishable and their entropies add extensively. However, once the particles approach sufficiently closely to interact via mutual forces, can they still be considered distinguishable?
@@
To address this, let us restrict the state space to a single conserved property: linear momentum in one dimension. The initially distinguishable particles now occupy two unique, distinguishable positions in this reduced state space. If a mutual force acts between them, for example, in an elastic collision, momentum is redistributed under conservation laws. After the interaction, the individual particle momenta no longer provide a reliable distinguishing feature. Thus, in the momentum degree of freedom, the initially distinguishable particles become indistinguishable after collision. Attempts to preserve distinguishability through continuous measurement only introduce further complications. While in principle classical properties of the two particles can be defined with arbitrary precision, this requires that the mutual forces driving momentum exchange also become arbitrarily negligible. In the limit of perfect precision, the particles are effectively non-interacting, as in the ideal gas model, which is not physically realistic. Moreover, defining distinguishability through measurement is problematic. Measurement must be performed using a probe particle. This probe must be initially at non-equilibrium with the ensemble, and ‘observes’ the microstates through thermalization, effectively performing a temperature measurement for each conserved quantity (see Appendix A). The probe particle (measurement device) necessarily interacts with the system during the thermalization with the system. For sufficiently small particles in the two particle collision, a probe would perturb the final momentum distribution. If the probe is non-perturbative or only weakly interacting, then insufficient information is obtained to guarantee distinguishability of the colliding particles [56].
@@
The difficulty of persistent distinguishability continues when additional conserved properties are reintroduced. If each conserved quantity is mutually independent, then they can only interact within their respective state-space axes, with the example already discussed now applying to each conserved quantity independently. From the information-theoretic viewpoint adopted in this paper, interactions correspond to the redistribution of information between particles, producing correlations that make particle states inseparable and hence indistinguishable. Even when conserved quantities are not mutually independent, interactions still generate mutual information that limits distinguishability in accordance with all conservation laws. Therefore, small classical particles cannot, even in principle, remain distinguishable from an entropic perspective if they interact. The fact that particles do interact, and that measured heat capacities are consistent with ensembles of indistinguishable particles [55], reinforces this conclusion.
@@
We are now in a position to propose a general description of particle indistinguishability from the perspective of information thermodynamics. What is fundamentally indistinguishable? Is it the conserved quantities themselves? Is it the particles? We argue that it is neither. Rather, it is information, or coherence in it’s most general form that is indistinguishable. The generalization of thermodynamics to arbitrary conserved quantities (Appendices AD), combined with the inseparability of probability distributions under interaction, supports this view. It also provides a natural intuition for the coherent transfer of information between mutually independent conserved quantities. It is precisely this redistribution of information, and the correlations it generates, that the entropy battery exploits.
Appendix G Endoreversible efficiency for two thermal reservoirs

### v2

@@ VII.1 Distinguishability
where $\mathbf{k}=\{k_{0},k_{1},...,k_{d-1}\}$ specifies a configuration of the $k_{j}$ particles over the $j$ states. See Appendix B.1 for further details. This partition function is immediately factorizable for every one of the $N$ particles, since it is identically a multinomial expansion
@@
where $p(\gamma)$ is the single particle Boltzmann distribution. The entropy at infinite temperature $\gamma=1/\tau=0$ is therefore proportional to the number of particles
@@
A separable probability distribution and extensive entropy can only occur if the particles are mutually independent. In the distinguishable particle ensemble, this mutual independence holds for each of the $N$ particles. Mutual independence therefore implies non-interaction, where interaction is defined generally as an event in which particles exchange a conserved quantity under the constraint of the relevant conservation law, or equivalently as a transfer of mutual information between particles. Consequently, distinguishable particles that maintain distinguishability at all times must be non-interacting, an argument similarly made by Jaynes in [26]. This has important implications for the entropy capacity of a realizable entropy battery. For unitary heat engines to operate, or for two canonical ensembles to thermalize, interactions between particles are required. For these reasons, distinguishable particles cannot be used in the entropy battery. Further discussion can be found in Appendix F.
The next best scenario is an ensemble of bosons. As discussed in Appendix B.2, unlike distinguishable particles, the probability distribution for an ensemble of bosons is inseparable over each of the particles. Therefore, as a contrasting argument to the distinguishable particles, bosons must allow interaction, such that one cannot describe the particles as individuals after there becomes an exchange of information. The description we give implies that it is the mutual information upon interaction that ultimately determines a particles distinguishability. If the information is mutually exclusive then there remains a form of distinguishability, yet no interaction, but if the information is mutually dependent then we lose distinguishability and there must have been an interaction. In this sense, what is ultimately indistinguishable upon interactions is the information that is encoded in the conserved quantities of the system. This is especially apparent when monitoring the mutual exchange of information between typically mutually independent conserved quantities like that used in the entropy battery. The exchange of generalised heat and work between conserved quantities is best understood under the context of information transfer.
@@
VIII Conclusion
From the perspective of information thermodynamics, an ensemble of particles with multiple, mutually independent conserved quantities exhibits an enlarged entropy capacity compared to classical, energy only canonical ensembles. By applying unitary operations that couple these conserved quantities, this entropy capacity can be harnessed during work extraction, yielding energy efficiencies beyond the Carnot limit. We refer to the ensemble as being an entropy battery, with it storing entropy as it thermalizes with an external reservoir. We justify treating spin angular momentum as an independent thermodynamic system by showing that, like energy, spin reservoirs obey the laws of generalized thermodynamics, follow Bose-Einstein and Fermi-Dirac statistics, satisfy the fluctuation-dissipation and equipartition theorems (see Appendix B), and have analogous Einstein solid and Debye models of heat capacity. We also discuss the conditions under which particles are distinguishable (Appendices B.1 and F), concluding that distinguishable particles can exist only if they are non-interacting. Consequently, the apparent entropy capacity of such particles cannot be used for energetic work extraction via heat engines. This leads naturally to a generalization of indistinguishability in informational terms, independent of the specific conserved quantities or particles, and provides a conceptual understanding of the entropy battery as a device that transfers entropy through subsystem entanglement.
@@ Acknowledgements
Data availability statement
The data that support the findings of this article are openly available [56].
@@ Appendix A Information Thermodynamics
In this section we derive, and present where assumed, the laws of spin thermodynamics, with these laws applying equally to arbitrary mutually independent conserved quantity.
@@
For ensembles with large $N$, the fluctuations of spin angular momentum and entropy between the baths become vanishingly small with respect to the average equilibrium values, as the relative variation scales with $1/\sqrt{N}$ [49]. We can therefore approximate the small changes as an infinitesimal, and make use of the result in eq.(33) to show that this can only be true if
@@ B.1.1 Polarization, spin temperature relation
It is convenient to re-express the spin temperature $\tau$ in terms of the spin-polarization $\alpha$, which provides a mapping to the macroscopic observable $J_{z}$. From eq.(35)
@@ B.2 Bosons
By changing the index from the particle configuration $\mathbf{k}$ to the macrostate $a(\mathbf{k})=\sum_{j=0}^{d-1}m_{j}k_{j}$, we show that spin bosons follow Bose-Einstein statistics [48]. The macrostate index $a(\mathbf{k})$ ranges from SN-SN to $SN$, which we can shift to the computational macrostate $m=a(\mathbf{k})+SN=\sum_{j=0}^{d-1}jk_{j}$ without affecting the distribution. The multiplicity using this index is derived in Appendix C, and is explicitly the coefficient of the $x^{m}]$th term of a Gaussian binomial,
@@
where $t^{N}]$ is a selection variable, used to select the polynomials of $t$ that correspond to $N$ particles, and eq.(48) is an identity from Newton’s generalized binomial theorem. Eq.(48) demonstrates that, in the infinite-particle limit, the boson partition function becomes separable over the macrostates
@@ B.3 Fermions
where $t^{N}x^{m}]$ selects the polynomial terms that correspondingly have $N$ atoms with macrostate index $m$. The fermion partition function using the $m$ index is then
@@
The fermion partition function is therefore separable over the macrostates $m$, allowing each state to be treated as a mutually independent grand canonical ensemble, where
@@
which is the Fermi-Dirac distribution [63]. We have removed the selection variable $t$, as this expression does not impose a constraint on the total particle number $N$. The separability of the probability distribution across the states $j$ allows each state to be treated as mutually independent. However, as in the bosonic case, the particles themselves cannot be considered mutually independent. Consequently, from a statistical perspective, fermions also permit interaction between particles. Finally, in the limit that $\tau\rightarrow\infty$, eq.(52) becomes
@@ Appendix C Multiplicity of Boson spin states
The number of configurations of indistinguishable bosons that yield the same total spin angular momentum can be reformulated as a combinatorial problem: counting the number of unique paths on an $N\times d$ grid, starting from the bottom-left corner and ending at the top-right corner, with movement restricted to upward and rightward steps, as illustrated in Fig.6. The grid represents $N$ particles distributed among $d$ states, and the area of the grid above the path taken represents the macrostate $m(\mathbf{k})=\sum_{j=0}^{d-1}jk_{j}$.
@@ Appendix D Waste and Entropic responses
as the amount of entropy the system can store per unit change in spin temperature. It should be clear that any mutually independent conserved quantity possesses an analogous heat capacity. It is therefore natural to define $C_{s}/\tau$ as the entropic response of the ensemble, a quantity that applies generally to any independent conserved variable, such as spin or energy. $C_{s}$ should be interpreted as the waste response of the ensemble, specific to the conserved quantity in question and associated with an analogous “heat” or the loss of usefulness of that conserved quantity. Re-naming $C_{s}$ from what is conventionally the ‘heat’ capacity emphasizes that $C_{s}$ is not a measure of capacity per se, but rather quantifies the absorbed ‘waste’ in response to a change in temperature. This is especially evident when considering the expression of the distinguishable waste response from table 2. The proportionality to the variance of the conserved quantity is a phenomena found generally in the fluctuation-dissipation theorem [49], where $C_{s}$ is a response function. Renaming the heat capacity to the waste response is therefore justified from this perspective. Eqs.(53) and (55) are presented as alternate, yet equivalent interpretations of a systems ability to store waste, both from a physical and an entropic perspective. These two equations allow for us to interpret heat engines in §IV as being machines that either filter conserved quantities by separating the ‘heat’ from the ‘work’, or machines that merely transport entropy and information between the heat baths, with these interpretation also applying to the operation of conventional energy heat engine. The information interpretation of the heat engine is general, and leads naturally to concepts involving information transfer between ‘waste’ baths of different conserved quantities, such as spin and energy.
@@
Finally, in the infinite $N$ limit the entropy and spin therm for bosons becomes eqs.(25) and (2) respectively. The boson spin therm is proportional to $\tau$ in the high temperature limit, as the Taylor expansion of $\coth(1/x)\approx x$ for $x\gg 0$, which re-derives the usual heat expression $\mathcal{Q}\propto\tau$ used in, for example, deriving the standard Carnot efficiency limit $\eta=1-\tau_{c}/\tau_{h}$. Eqs.(25) and (2) are presented graphically in Fig.8.
@@ E.1 Distinguishable particles
From eq.(39), the partition function for an ensemble of $N$ distinguishable particles with $d$ states per particle is
@@
for probability $P(j,\tau)=e^{-j/\tau}/Z$, and $Z=\sum_{j=0}^{d-1}e^{-j/\tau}$. Note also that $\braket{j}N=\braket{\bar{J}_{z}}+SN$. The second derivative is
@@
$\frac{C_{s}}{N}=\frac{1}{\tau^{2}}\text{Var}(\bar{J}_{z}(\tau)).$
@@ E.2 Bosons
From appendix B.2, the partition function for an ensemble of $N$ boson with $d$ states per particle is
@@ E.3 Debye and Einstein models
This is for 3 degrees of freedom with $N$ particles, energy $h\nu$, and units J/K thanks to the presence of the Boltzmann constant. For 1 degree of freedom, the unitless, per particle heat capacity is
@@
where we have made the equivalent substitution in the spin degree of freedom, with $2S$ typically being the Debye cut off frequency $\nu_{D}$. Eq.(61) is equal to the boson model (60). The Debye model approximates the sum using an integral, while including the density of frequency modes [50], which is $j^{2}dj$ in the spin degree of freedom, resulting in the un-normalized Debye model
@@ Appendix F Distinguishability discussion
With the goal of this article being to present a realizable entropy storage device, and with entropy capacity translating into extractable energetic work as demonstrated by the entropy battery, then we require an understanding of how one could increase entropy capacity by means other than enlarging the size of the system. As already discussed in §VII and appendix B.1, aside from increasing the number of particles, number of states or the number of independent conserved quantities, the only other apparent means of increasing entropy capacity is to render each particle distinguishable. However, as also established in those sections, a system of distinguishable particles must necessarily be non-interacting. Since particles do in fact interact and heat capacities for distinguishable particles are very different from experimental heat capacities, then particles must be indistinguishable. In this section, we examine whether it is possible to retain the physical entropy capacity associated with distinguishable particles while still allowing interaction. Our conclusion, inspired largely by Saunders’ analysis of indistinguishability [25], is that this is not possible even in principle. The reader is referred to his work for a thorough treatment of the indistinguishability problem.
@@
We begin by clarifying what we mean by distinguishability. From an entropic perspective, physical distinguishability increases the number of accessible microstates $\Omega$ available to the ensemble, leading to an enlarged entropy capacity. Therefore, distinguishability is real when the entropy capacity of the ensemble at a temperature is larger than that of an ensemble of indistinguishable particles at the same temperature. The entropy capacity is determined by the total number of microstates, which in itself is determined by the number of unique distributions of the particles in state space. Under this definition, if two ensemble states are physically identifiable, then these are distinct microstates, with each contributing to the system’s physical entropy and hence to its extractable work.
@@
Consider first a single classical particle with physical properties such as position, momentum, and charge. Let these properties be localised in state space (a minimum entropy state). Since there is only a single particle, then the particle can be tracked continuously through unitary evolution, rendering it distinguishable at all times. Now introduce a second identical particle (same physical properties), however occupying a different location in state space. As long as these particles do not interact and their probability distributions remain separable, they are distinguishable and their entropies add extensively. However, once the particles approach sufficiently closely to interact via mutual forces, can they still be considered distinguishable?
@@
To address this, let us restrict the state space to a single conserved property: linear momentum in one dimension. If a mutual force acts between them, for example, in an elastic collision, momentum is redistributed under the law of conservation of linear momentum. After the interaction, due to this redistribution, the individual particles momenta no longer act as a reliably distinguishable feature. Therefore, in the momentum degree of freedom, the initially distinguishable particles become indistinguishable after collision. Attempts to preserve distinguishability through continuous measurement only introduce further complications. Consider what happens when the two particles are classical. In principle, their properties can be defined with arbitrary precision. However, this implies that the mutual forces responsible for momentum exchange must become arbitrarily small. If any interaction were present, information would be shared between particles, making their individual states interdependent and less distinguishable. In other words, the act of interacting necessarily introduces correlations that blur the precise definition of each particle’s state. Therefore, in the limit of perfect precision, the particles must be effectively non-interacting, just as assumed in the ideal gas model. Moreover, defining distinguishability through measurement is problematic, as measurement must be performed using a probe particle, or via some interaction with the system of interest. The probe must initially be in a known low entropy state, which is at non-equilibrium with the ensemble, and the probe ‘observes’ the microstates of the ensemble through an effective thermalization. The measurement is in effect a temperature measurement for each conserved quantity (see Appendix A). The probe particle (measurement device) necessarily interacts with the system during the effective thermalization. If the two particles that we intend on distinguishing are sufficiently small particles, the probe will perturb the final momentum distribution, removing the possibility of distinguishing the particles in a subsequent measurement, and ruling out continuous measurement as a means for enabling distinguishability. If the probe is non-perturbative or only weakly interacting, then insufficient information is obtained to guarantee distinguishability of the colliding particles [68].
@@
The difficulty of persistent distinguishability during the two particle collision continues when additional conserved properties are reintroduced. If each conserved quantity is mutually independent, then interactions only transfer information within their respective state-space planes, with the example of linear momentum in the prior paragraph now applying to each independent conserved quantity. Even when conserved quantities are not mutually independent, interactions still generate mutual information that limits distinguishability in accordance with all conservation laws. Therefore, small classical particles cannot remain distinguishable from an entropic perspective if they interact, even in principle. The fact that particles do interact, and that measured heat capacities are consistent with ensembles of indistinguishable particles [25], reinforces this conclusion.
@@
We are now in a position to propose a general description of particle indistinguishability from the perspective of information thermodynamics. What is it that makes particles fundamentally indistinguishable? Is it the conserved quantities themselves? Is it the particles? We argue that it is neither. Rather, it is information, or coherence in it’s most general form that is indistinguishable. This assessment is in agreement with the indistinguishability of information in quantum mechanics, although argued from the perspective of classical mechanics. The generalization of thermodynamics to arbitrary conserved quantities (Appendices AD), combined with the inseparability of probability distributions after interaction, supports this view. It also provides a natural intuition for the coherent transfer of information between mutually independent conserved quantities. It is precisely this redistribution of information that the entropy battery exploits.
Appendix G Endoreversible efficiency of the entropy battery